DPI Logo

DPI Logo

Friday, January 27, 2017

ELO Problems Today 1/27/2017

Our ELO vendor did a planned server upgrade last night, and there were unexpected results this morning.  Educators could not complete their applications, and preparation programs could not upload ECL data.

We had to take the system down around lunch so that the vendor could make fixes and restart servers.  We believe that the system is now working properly and hope this was not a great inconvenience.

No comments:

Post a Comment